# Dynamical systems approach to Cold and Warm Inflation within slow-roll and beyond

### v1

@@ Abstract
In this work, we systematically present a new dynamical systems approach to standard inflationary processes and their variants as constant-roll inflation. Using the techniques presented in our work one can in general investigate the attractor nature of the inflationary models in the phase space. We have compactified the phase space coordinates, wherever necessary, and regulated the nonlinear differential equations, constituting the autonomous system of equations defining the dynamical system, at the cost of a new redefined time variable which is a monotonic increasing function of the standard time coordinate. We have shown that in most of the relevant cases the program is executable although the two time coordinates may show different durations of cosmological events. If one wishes one can revert back to the cosmological time via an inverse transformation. The present work establishes a standard norm for studying dynamical as well as stability issues in any new inflationary system.
@@ 1 Introduction
In general the stability issues about any system are best studied in the dynamical systems approach where one recasts all the dynamical equations of the system in the form of nonlinear, first order autonomous system of differential equations. The dynamical systems approach has emerged as a powerful tool in late-time cosmology, particularly for analyzing new classes of cosmological models [13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31]. In inflationary cosmology, one rarely use the full potential of the dynamical systems approach as we do not expect the system to have any stable fixed points during the inflationary regime. Moreover, a successful period of inflation ends in reheating the universe and the physics of reheating [32, 33, 34, 35] is different from the physics of slow-roll inflation. Consequently a dynamical system which models a slow-roll process will be unable to yield meaningful information near the end of inflation unless the equations are modified so that the reheating process is also included in the dynamical equations. It is practically impossible to describe the initiation of inflation and the reheating phase with the same set of dynamical equations and mostly the dynamical systems approach concentrates on the initial development of an inflationary system and the analysis can be extended approximately up to the graceful exit phase. In warm inflation (WI) one does not have a separate reheating phase and we expect that a dynamical systems approach in WI is more useful. A dynamical systems approach has some advantages, it is a very general method which has the potential to unravel the dynamical properties of any system unambiguously. If the dynamical systems approach can be used in conjunction with a wide class of initial conditions then we get a set of trajectories in the phase space showing the qualitative behavior of the system. Except giving an overall, qualitative phase space picture of inflationary dynamics one can also analyze any inflationary scenario in detail using the methods of dynamical systems. In this paper we have tried to posit a uniform dynamical systems viewpoint using which one can study any inflationary process.
@@
This last condition is essential for producing thermal fluctuations. Few previous authors have studied the dynamical systems approach and stability issues in WI [53, 54, 55, 56, 57, 58, 59, 60]. In the present work we generalize the previous techniques used by these authors and try to give new directions for the qualitative phase space study of WI. We have shown the attractor nature of WI where the dynamical system is highly constrained by the dynamical equilibrium requirements of the radiation bath. In general if one is purely interested in WI dynamics for a particular initial condition, then one may not use the constraints imposed by the equilibrium of the radiation bath because in slow-roll regime the variation of the temperature of the bath is relatively small during the inflationary period. In our case we produce the phase space trajectories of the system for various possible initial conditions: all of these trajectories may not be undergoing a slow-roll process. To differentiate the phase space trajectories which evolve through a thermalized phase and non-thermalized phase we have explicitly shown how one can formulate the constrained WI dynamical system.
Closely related to the above two main modern paradigms of inflationary dynamics are some new variants. These variants are all some forms of CI or WI but they differ from the canonical models by the scalar field rolling condition. One of these variants are constant-roll cold inflation (CRCI) [61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77], which is a variant of CI where the slow-roll (SR) condition of the scalar field is changed to the constant-roll condition. In general inflationary dynamics is specified by some rolling conditions and one of these conditions are related to the value of $\ddot{\phi}/(H\dot{\phi})$ where $\phi$ represents the inflaton field and $H$ is the Hubble parameter during inflation. The dot represent derivatives with respect to cosmological time. In the slow-roll regime this ratio tends to zero, whereas in the constant-roll regime this ratio remains a constant. The constant-roll (CR) condition restricts the inflationary system although it is known that it remains an attractor solution near the SR limit. In this work we will verify this claim and show how the attractor solution in CRCI gets modified from the attractor solution in CI. We will also show that in CRCI there will be some dynamical evolutions where an accelerated expansion phase, like the inflation phase, goes on inflating and cannot come out of the inflationary process through graceful exit.
@@
In this paper, we have independently presented the dynamical analysis for each type of inflation discussed, allowing for a direct comparison of how different inflationary conditions influence the phase space dynamics. For the first time, a unified framework based on dynamical systems theory has been employed to study the various forms of inflation. We have successfully applied compactified phase space techniques and addressed ill-defined autonomous equations through a redefinition of time—methods that are novel in the context of inflationary dynamics. An exception is made in the case of CRCI, where the dynamical phase space is effectively one-dimensional. In such cases, a conventional autonomous system analysis offers limited insight. Therefore, for CRCI, we have adopted an alternative approach using stream plots of the phase space variables in a two-dimensional plane.
The material in this paper is organized as follows. In the next section 2, we briefly introduce the fundamentals of dynamical systems theory and apply it to analyze the phase space dynamics in cold inflation (CI). In section 3, we discuss the dynamics of CRCI for various values of the constant-roll parameter. Section 4 presents the phase space analysis for warm inflation (WI), wherein we compactify the phase space and regulate the dynamical system through a redefinition of time. In section 5, we extend the dynamical systems analysis to the constant-roll warm inflation (CRWI) case for different values of the constant-roll parameter. Finally, we summarize our approach and highlight the key findings regarding inflationary dynamics in the concluding section 6.
2 A brief introduction to the theory of dynamical systems and its application to the standard slow-roll cold inflation model
In this section we specify the conventions and notations used in this paper. Initially we specify the mathematical principles, related to the dynamical systems approach, which are used to analyze the various inflationary cases.
@@
where $t$ is the cosmic time, $a(t)$ is the scale-factor of the universe, and $\mathbf{x}$ represents the spatial coordinates $(x,y,z)$. The fundamental dynamical equations governing a cold inflationary system consist of the Friedmann equations and the Klein–Gordon equation for the inflaton field, formulated in the FLRW background. In the case of warm inflation, the presence of a subdominant yet non-negligible radiation fluid—coupled to the inflaton field—necessitates the inclusion of the radiation fluid’s evolution equation as part of the dynamical system. To apply the dynamical systems approach, it is customary to recast the equations into an autonomous system using suitably defined phase space variables. These variables are chosen to ensure that the phase space has the minimal possible dimensionality and that all quantities involved are dimensionless. In our analysis, we reparametrize the cosmic time $t$ either using a dimensionless time variable $t^{\prime}$ in certain cases, or by the number of $e$-folds, defined via $dN=H\,dt$, where $H\equiv\dot{a}/a$ is the Hubble parameter.
@@
2.1 Dynamical analysis of the standard cold inflation in slow-roll regime
Here, to set the stage, we will analyze the dynamics of a simplest slow-roll cold inflation model in the phase space using the techniques mentioned above. In standard cold inflation, the dynamics of the inflaton field $\phi$ is governed by the Klein-Gordon equation
@@ 2.1 Dynamical analysis of the standard cold inflation in slow-roll regime
where $V(\phi)$ is the inflaton potential and the subscript $\phi$ indicates partial derivatives with respect to $\phi$. Assuming the Universe is dominated by the inflaton field during cold inflation, the Friedmann equations take the form:
@@
where $\kappa^{2}=1/M_{\rm Pl}^{2}$, and the reduced Planck mass is defined as $M_{\rm Pl}=1/\sqrt{8\pi G}$ in natural units ($c=\hbar=1$), with $G$ denoting Newton’s gravitational constant. The slow-roll parameter $\epsilon_{1}$, defined as
@@
determine the equation of state $\omega_{\phi}$ of the inflaton field as
@@
showing that during a slow-roll phase, when $\epsilon_{1}\ll 1$, $\omega_{\phi}$ tends to $-1$.
To demonstrate the phase space dynamics in this scenario, we choose the quadratic self potential for the inflaton field as
$V(\phi)=\frac{1}{2}m^{2}\phi^{2},$ (2.11)
where $m$ is the mass of the inflaton field. While the dynamics of such a system have been extensively studied in the literature, we present a brief review for continuity. We define the dimensionless dynamical variables in the following way
$x=\pm\frac{\kappa\sqrt{V}}{\sqrt{3}m},\quad y=\frac{\kappa\dot{\phi}}{\sqrt{6}m},\quad z=\frac{H}{m},$ (2.12)
where the Hubble parameter $H$, potential $V$, and the field velocity $\dot{\phi}$ are normalized with respect to $m$. The Hubble constraint relation using the Friedmann equation given in Eq. (2.6) can be expressed in terms of the predefined variables as:
$z^{2}=y^{2}+x^{2}\ .$ (2.13)
The fractional energy density parameter $\Omega_{\phi}$, the equation of state (E0S) $\omega_{\phi}$, and $\epsilon_{1}$ can be expressed in terms of these dimensionless dynamical variables as
$\Omega_{\phi}=\frac{\kappa^{2}\rho_{\phi}}{3H^{2}}=\frac{x^{2}+y^{2}}{z^{2}}=1,\quad\omega_{\phi}=\frac{y^{2}-x^{2}}{y^{2}+x^{2}},\quad\epsilon_{1}=\frac{3y^{2}}{z^{2}}=\frac{3y^{2}}{x^{2}+y^{2}}\,.$ (2.14)
It is to note that defining the dynamical variables as above makes them vary as
$-\infty<x<\infty,\quad-\infty<y<\infty,\quad 0<z<\infty\,,$ (2.15)
yielding the phase space to be noncompact. This noncompact phase space can be mapped into a compact space using the Poincaré transformation [81] as:
$\bar{x}=\frac{x}{\sqrt{1+x^{2}}},\quad\bar{z}=\frac{z}{\sqrt{1+z^{2}}}\,.$ (2.16)
Looking at the inverse transformations,
$x=\frac{\bar{x}}{\sqrt{1-\bar{x}^{2}}},\quad z=\frac{\bar{z}}{\sqrt{1-\bar{z}^{2}}}\,,$ (2.17)
it is evident that the transformed variables $(\bar{x},\bar{z})$ vary over finite ranges:
$-1<\bar{x}<1,\quad 0<\bar{z}<1\,.$ (2.18)
We can further compactify $y$ by a similar redefinition,
$\bar{y}=\frac{y}{\sqrt{1+y^{2}}}\ .$ (2.19)
ensuring that $\bar{y}$ varies over a finite range $-1<\bar{y}<1$ as well. With this compactification, the variable $y$ is effectively constrained to a bounded interval, making the phase space more manageable for analysis and visualization. Substituting this transformation into the constraint condition leads to:
$\frac{\bar{z}^{2}}{{1-\bar{z}^{2}}}=\frac{\bar{x}^{2}}{{1-\bar{x}^{2}}}+\frac{\bar{y}^{2}}{1-\bar{y}^{2}}\ .$ (2.20)
This compactified form ensures that all variables ($\bar{x}$, $\bar{y}$, and $\bar{z}$) are finite and their dynamics can be studied within a well-defined and bounded phase space.
Figure 1: The phase space of inflaton field in the variable $(\bar{x},\bar{y})$, where yellow region represents $0.10<\epsilon_{1}<$ 0.1, green region shows $10.1<\epsilon_{1}<$ 1 and pink region signifies $\epsilon_{1}>1$, and .
The Klein-Gordon equation (Eq. (2.5)) and the second Friedmann equation (Eq. (2.7)) can be recast in terms of the dynamical variables $\bar{x}$ and $\bar{y}$ which constitute the autonomous equations of this system:
$\bar{x}^{\prime}$ $=$ $\frac{\bar{y}\left(1-\bar{x}^{2}\right)^{3/2}}{\sqrt{1-\bar{y}^{2}}}\,,$ (2.21)
$\bar{y}^{\prime}$ $=$ $\frac{3\bar{y}\sqrt{1-\bar{y}^{2}\bar{x}^{2}}\sqrt{\left(1-2\bar{y}^{2}\right)\bar{x}^{2}+\bar{y}^{2}}}{\left(1-\bar{y}^{2}\right)^{3/2}\left(\bar{x}^{2}-1\right)}-\frac{\bar{x}}{\sqrt{1-\bar{x}^{2}}}\ ,$ (2.22)
where the prime denotes derivatives with respect to the dimensionless time $t^{\prime}=mt$. It is to note that now the dynamical equations are written in the form given in Eqs. (2.3) and have no further pathologies. It is then straightforward to write down the Jacobian as in Eq. (2.4) and find the fixed points. The analysis yields a single critical point at $P_{1}=(0,0)$, where both $\bar{x}$ and $\bar{y}$ vanish. At $P_{1}$ both $\phi$ and $\dot{\phi}$ vanishes and consequently the Hubble parameter is also zero. It turns out to be a stable fixed point. It must be stated here that $P_{1}$ is the only fixed point in this case, and there are no fixed points at infinity as in the compact phase space the points at infinity have been mapped to the borders $\bar{x}=\pm 1$ and $\bar{y}=\pm 1$. Any possible critical points at infinity would have appeared as critical points on those lines. Fig. [1] illustrates the phase-space trajectories, with color-coded regions: the yellow and green regions correspond to spaces where $0.10<\epsilon_{1}<$ 0.1 and $10.1<\epsilon_{1}<$ 1 and hence these regions are the slow-roll regions, whereas the pink region corresponds to portions of the phase space where $\epsilon_{1}>1$. On the boundary curve between the pink region and the green region (bordered in red) we have $\epsilon_{1}=1$.
As far as cold inflation is concerned, we know that in such inflationary cases the slow-roll parameter starts with a value much less than unity and inflation goes on until $\epsilon_{1}\sim 1$. Consequently in the phase space of the system inflationary trajectories are those which start in the yellow region and continuously travels towards the pink region through the green region. We can spot such inflationary trajectories both on the left and right sides of the fixed point $P_{1}$ in Fig. [1]. Here we must remind the reader that some of these trajectories may seem to again enter a inflationary phase as shown in the phase space plot. Does the phase space imply multiple inflationary phases? The answer is no. The reason being the plot in Fig. [1] is purely a dynamical systems analysis of the cold inflationary model and does not contain any information about reheating and graceful exit. In reality as soon as one of the trajectories starts from the yellow region and nears the pink region the physics of the system changes as graceful exit takes place. The phase space plot only gives us the information about those trajectories that may produce inflation.
The phase space picture shows that most of the trajectories, corresponding to various kinds of initial conditions, ends up in an inflationary phase. In this sense cold inflation is an attractor solution. Although we have shown the attractor like solution with respect to a particular inflaton potential, for any other potential which allows inflation will have a similar kind of phase space structure.
3 Dynamical analysis of cold inflation in the constant-roll regime
@@
Constant-roll cold inflation (CRCI) was first introduced in [83, 84], and the theory of CRCI was subsequently expanded and developed in [62, 64, 65, 63]. Some of these papers have also carried out a dynamical system analysis of the inflationary process in constant-roll regime [66, 64]. However, these previous studies were not solely dedicated to fully analyze the attractor nature of CRCI in a detailed manner. Here, we develop the general techniques and explore the dynamical system of CRCI. In CRCI all the essential dynamical equations of cold inflation, as discussed in the previous subsection, are valid in addition with an extra constant-roll condition given by:
@@ 3.1 Case I: $\mathbf{\beta>0}$
which can always be fulfilled if $C\sim\pi/2$ and adjusting the value of $M$. This condition specifies $H>0$. As the scale-factor $a$ is always assumed to be greater than zero, implies $\dot{a}>0$. The authors of the above mentioned paper opines (after Eq. (2.29) of their work) that “Although this is a mathematically allowed solution, it has $\ddot{a}(t)<0$. Therefore, it cannot describe an inflationary model in the usual sense.” We show that such a conclusion is erroneous, there exists initial conditions and parameter values for which both $\dot{a}>0,\ \ \ddot{a}>0$.
@@
In the limit $3\beta\to 0$, we see that the above expression can indeed produce $\ddot{a}>0$ during some inflationary epoch. Only when $13\beta>$ 1 we do not find any possibility of inflation.
To analyze the dynamics associated with the potential given in Eq. (LABEL:cr_v_betac_positive), we begin with the equation of motion for the inflaton field $\phi$ as given in Eq. (2.5). Substituting the constant-roll condition from Eq. (3.1), the modified equation of motion for the inflaton field becomes:
$3(1-\beta)H\dot{\phi}+V_{,\phi}=0$ (3.9)
@@
These dynamical variables are structurally similar to those defined for the earlier case and yield the same value of the slow-roll parameter used in the earlier case: $\epsilon_{1}=3y^{2}/(x^{2}+y^{2})$. Due to the different constants involved in the potential, we normalize the field and the Hubble parameter with the constant $V_{0}$, thereby constructing dimensionless dynamical variables. One must note that in the present case the potential in Eq. (LABEL:cr_v_betac_positive) is not positive definite. This potential may admit negative values for some parameter choice and some values of the initial conditions. As because negative value of the scalar field potential can never produce inflation, we have worked with those initial conditions for which $V(\phi)$ is non-negative and consequently $x$ is properly defined everywhere in the phase space. From the form of the potential it is evident that when $\phi\sim\phi_{0}$, $V(\phi)$ turns negative and consequently in the present case we only analyze those dynamical regions where $x>0$ or $x<0$ as $x$ never crosses zero.
@@
Figure 2: The streamline plots for the field $(\kappa\phi,\dot{\phi}/\sqrt{V_{0}})$ in a specific $(x,y)$ range. The curves correspond to streamlines for different values of $\beta>0$ and field constants. The color code is as follows: the green region represents regions where $10<\epsilon_{1}<$ 1, in the pink region $\epsilon_{1}>1$ and in the yellow region we have $0.050<\epsilon_{1}<$ 0.05.
which gives $z$ in terms of $x$ and $y$. We have produced the streamline plots for the field $(\dot{\phi}/\sqrt{V_{0}},\kappa\phi)$ in some specific range of $x$ and $y$. The stream lines are the flow lines for the vector field $(\kappa\phi,\dot{\phi}/\sqrt{V_{0}})$. The stream-plots are shown in Fig. [2] for different values of $\beta$ and $\kappa\phi_{0}$. Each stream-plot is divided into distinct colored regions: the green region represents $10<\epsilon_{1}<$ 1, the pink region represents $\epsilon_{1}>1$, and the yellow region represents $0.050<\epsilon_{1}<$ 0.05. On the boundary curve, differentiating the green and the pink regions, $\epsilon_{1}=1$.

### v3

# Dynamical Systems Approach to Non-Slow-Roll Inflationary Models

@@ Abstract
In this work, we systematically present a new dynamical systems approach to nonstandard inflationary processes as constant-roll inflation and ultraslow-roll inflation. Using the techniques presented in our work one can in general investigate the attractor nature of the inflationary models in the phase space. We have compactified the phase space coordinates, wherever necessary, and regulated the nonlinear differential equations, constituting the autonomous system of equations defining the dynamical system, at the cost of a new redefined time variable which is a monotonic increasing function of the standard time coordinate. We have shown that in most of the relevant cases the program is executable although the two time coordinates may show different durations of cosmological events. Our methods of analysis differs slightly in different models but we have always emphasized on the nature of the initial conditions leading to stable inflationary phases in different cases. We have provided a universal language in terms of which various nonstandard inflationary models can be studied.
@@ 1 Introduction
In general the stability issues about any system are best studied in the dynamical systems approach where one recasts all the dynamical equations of the system in the form of nonlinear, first order autonomous system of differential equations. The dynamical systems approach has emerged as a powerful tool in late-time cosmology, particularly for analyzing new classes of cosmological models [24, 23, 6, 25, 89, 31, 83, 37, 22, 39, 28, 47, 48, 49, 19, 32, 85, 76]. In inflationary cosmology, one rarely use the full potential of the dynamical systems approach as we do not expect the system to have any stable fixed points during the inflationary regime. Moreover, a successful period of inflation ends in reheating the universe and the physics of reheating [56, 57, 3, 8] is different from the physics of slow-roll inflation. Consequently a dynamical system which models a slow-roll process will be unable to yield meaningful information near the end of inflation unless the equations are modified so that the reheating process is also included in the dynamical equations. It is practically impossible to describe the initiation of inflation and the reheating phase with the same set of dynamical equations and mostly the dynamical systems approach concentrates on the initial development of an inflationary system and the analysis can be extended approximately up to the graceful exit phase. In warm inflation (WI) one does not have a separate reheating phase and we expect that a dynamical systems approach in WI is more useful. A dynamical systems approach has some advantages, it is a very general method which has the potential to unravel the dynamical properties of any system unambiguously. If the dynamical systems approach can be used in conjunction with a wide class of initial conditions then we get a set of trajectories in the phase space showing the qualitative behavior of the system. Except giving an overall, qualitative phase space picture of inflationary dynamics one can also analyze any inflationary scenario in detail using the methods of dynamical systems. In this paper we have tried to posit a uniform dynamical systems viewpoint using which one can study any inflationary process.
@@
This last condition is essential for producing thermal fluctuations. Few previous authors have studied the dynamical systems approach and stability issues in WI [35, 68, 58, 34, 79, 33, 93, 51].
Closely related to the above two main modern paradigms of inflationary dynamics are some new variants. These variants are all some forms of CI or WI but they differ from the canonical models by the scalar field rolling condition. One of these variants is the constant-roll cold inflation (CRCI) [70, 92, 43, 65, 88, 60, 72, 71, 77, 80, 40, 50, 54, 30, 4, 41, 42, 75, 5, 78], which is a variant of CI where the slow-roll (SR) condition of the scalar field is changed to the constant-roll condition. In general inflationary dynamics is specified by some rolling conditions and one of these conditions is related to the value of $\ddot{\phi}/(H\dot{\phi})$ where $\phi$ represents the inflaton field and $H$ is the Hubble parameter during inflation. The dot represent derivatives with respect to cosmological time. In the slow-roll regime this ratio tends to zero, whereas in the constant-roll regime this ratio remains a constant. The constant-roll (CR) condition restricts the inflationary system although it is known that it remains an attractor solution near the SR limit. In this work we will verify this claim and show how the attractor solution in CRCI gets modified from the attractor solution in CI. We will also show that in CRCI there will be some dynamical evolutions where an accelerated expansion phase, like the inflation phase, goes on inflating and cannot come out of the inflationary process through graceful exit.
@@
Another variant of the standard SR inflation is the ultraslow-roll (USR) inflation, which happens when $\ddot{\phi}/(H\dot{\phi})\sim-1$ when $V_{,\phi}\sim 0$, where $V_{,\phi}$ is the derivative of inflaton potential with respect to the inflaton field. In the present paper we have worked in warm USR model [20] where the USR condition is modified. It is well known that USR phase is a transient phase which terminates after a few $e$-folds and after that the SR phase commences. The outcome of our analysis is very interesting in this case. For a particular potential we can show that warm USR inflation does not have a general attractor like solution in the sense that, if the initial conditions do not satisfy the USR condition then USR will not lead to a SR phase. This condition is different from the previous cases where the system was always in a CR regime. CRCI or CRWI are not transient phases, whereas USR is a transient phase which terminates soon after its onset. In this case one may try to see whether more general initial conditions can naturally give rise to an USR phase.
In this paper, we have independently presented the dynamical analysis for each type of inflation discussed, allowing for a direct comparison of how different inflationary conditions influence the phase space dynamics. A unified framework based on dynamical systems theory has been employed to study the various forms of inflation. We have successfully applied compactified phase space techniques and addressed ill-defined autonomous equations through a redefinition of time. An exception is made in the case of CRCI, where the dynamical phase space is effectively one-dimensional. In such cases, a conventional autonomous system analysis offers limited insight. Therefore, for CRCI, we have adopted an alternative approach using stream plots of the phase space variables in a two-dimensional plane.
@@
where $t$ is the cosmic time, $a(t)$ is the scale-factor of the universe, and $\mathbf{x}$ represents the spatial coordinates $(x,y,z)$. The fundamental dynamical equations governing a cold inflationary system consist of the Friedmann equations and the Klein–Gordon equation for the inflaton field, formulated in the FLRW background. In the case of warm inflation, the presence of a subdominant yet non-negligible radiation fluid—coupled to the inflaton field—necessitates the inclusion of the radiation fluid’s evolution equation as part of the dynamical system. To apply the dynamical systems approach, it is customary to recast the equations into an autonomous system using suitably defined phase space variables. These variables are chosen to ensure that the phase space has the minimal possible dimensionality and that all quantities involved are dimensionless. In our analysis, we reparametrize the cosmic time $t$ either using a dimensionless time variable $t^{\prime}$ in certain cases, or by the number of $e$-folds, defined via dN=HdtdN=H\,dt, where $H\equiv\dot{a}/a$ is the Hubble parameter.
@@
The material in this paper is organized as follows. In the next section 2, we discuss the dynamics of CRCI for various values of the constant-roll parameter. Section 3 presents the phase space analysis for constant-roll warm inflation (CRWI), wherein we compactify the phase space and regulate the dynamical system through a redefinition of time. In section 4, we extend the dynamical systems analysis to the case of warm ultraslow-roll inflation. Finally, we summarize our approach and highlight the key findings regarding inflationary dynamics in the concluding section 5.
2 Dynamical analysis of cold inflation in the constant-roll regime
In standard cold inflation, the dynamics of the inflaton field $\phi$ is governed by the Klein-Gordon equation
@@ 2 Dynamical analysis of cold inflation in the constant-roll regime
where $V(\phi)$ is the inflaton potential and the subscript $\phi$ indicates partial derivatives with respect to $\phi$. Assuming the Universe is dominated by the inflaton field during cold inflation, the Friedmann equations take the form
@@
where $\kappa^{2}=1/M_{\rm Pl}^{2}$, the reduced Planck mass is given by $M_{\rm Pl}=1/\sqrt{8\pi G}$ where $G$ is the Newtonian gravitational constant. The Hubble slow-roll parameter slow-roll parameter $\epsilon_{1}$, defined as
@@
determine the equation of state $\omega_{\phi}$ of the inflaton fluid as
@@
This shows that during a slow-roll phase, when $\epsilon_{1}\ll 1$, $\omega_{\phi}$ tends to 1-1.
@@
Constant-roll cold inflation (CRCI) was first introduced in [62, 70], and the theory of CRCI was subsequently expanded and developed in [92, 65, 88, 43]. Some of these papers have also carried out a dynamical system analysis of the inflationary process in constant-roll regime [60, 65]. However, these previous studies were not solely dedicated to fully analyze the attractor nature of CRCI in a detailed manner. Here, we develop the general techniques and explore the dynamical system of CRCI. In CRCI all the essential dynamical equations of cold inflation are valid in addition with an extra constant-roll condition given by:
@@ 2.1 Case I: $\mathbf{\beta>0}$
which can always be fulfilled if $C\sim\pi/2$ and adjusting the value of $M$. This condition specifies $H>0$. Here, the constant $C$ is a physically meaningful parameter emerging from the constant-roll condition. Different choices of $C$ lead to different cosmological evolutions: for $C=0$, the Hubble parameter is negative near $t=0$, ruling out inflation, whereas for other values of $C$, one obtains $H>0$ near $t=0$, enabling an inflationary phase. In this sense, $C$ effectively encodes a freedom in redefining the origin of time. As the scale-factor $a$ is always assumed to be greater than zero, implies $\dot{a}>0$. The authors of the above mentioned paper opines (after Eq. (2.29) of their work) that “Although this is a mathematically allowed solution, it has $\ddot{a}(t)<0$. Therefore, it cannot describe an inflationary model in the usual sense.” We show that such a conclusion is erroneous, there exists initial conditions and parameter values for which both $\dot{a}>0,\ \ \ddot{a}>0$.
@@
In the limit $3\beta\to 0$, we see that the above expression can indeed produce $\ddot{a}>0$ during some inflationary epoch. Only when $3\beta>1$ we do not find any possibility of inflation.
To analyze the dynamics associated with the potential given in Eq. (2.9), we begin with the equation of motion for the inflaton field $\phi$ as given in Eq. (2.1). Substituting the constant-roll condition from Eq. (2.7), the modified equation of motion for the inflaton field becomes:
$3(1-\beta)H\dot{\phi}+V_{,\phi}=0.$ (2.15)
@@
These dynamical variables are structurally similar to those defined for the earlier case and yield the same value of the slow-roll parameter used in the earlier case: $\epsilon_{1}=3y^{2}/(x^{2}+y^{2})$. Due to the different constants involved in the potential, we normalize the field and the Hubble parameter with the constant $V_{0}$, thereby constructing dimensionless dynamical variables. One must note that in the present case the potential in Eq. (2.9) is not positive definite. This potential may admit negative values for some parameter choice and some values of the initial conditions. As because negative value of the scalar field potential can never produce inflation, we have worked with those initial conditions for which $V(\phi)$ is non-negative and consequently $x$ is properly defined everywhere in the phase space. From the form of the potential it is evident that when $\phi\sim\phi_{0}$, $V(\phi)$ turns negative and consequently in the present case we only analyze those dynamical regions where $x>0$ or $x<0$ as $x$ never crosses zero.
@@
Figure 1: The streamline plots for the field $(\kappa\phi,\dot{\phi}/\sqrt{V_{0}})$ in a specific $(x,y)$ range. The curves correspond to streamlines for different values of $\beta>0$ and field constants. The color code is as follows: the green region represents regions where $0<\epsilon_{1}<1$, in the pink region $\epsilon_{1}>1$ and in the yellow region we have $0<\epsilon_{1}<0.05$.
which gives $z$ in terms of $x$ and $y$. We have produced the streamline plots for the field $(\dot{\phi}/\sqrt{V_{0}},\kappa\phi)$ in some specific range of $x$ and $y$. The stream lines are the flow lines for the vector field $(\kappa\phi,\dot{\phi}/\sqrt{V_{0}})$. The stream-plots are shown in Fig. [1] for different values of $\beta$ and $\kappa\phi_{0}$. Each stream-plot is divided into distinct colored regions: the green region represents $0<\epsilon_{1}<1$, the pink region represents $\epsilon_{1}>1$, and the yellow region represents $0<\epsilon_{1}<0.05$. On the boundary curve, differentiating the green and the pink regions, $\epsilon_{1}=1$.